In how many ways can 10 instructors be assigned to eight sections of a course in mathematics?
Answer by jim_thompson5910(35256)

Assuming that order matters, there are 10 P 8 = (10!)/(10-8)! = 10!/2! = 3628800/2 = 1814400 different ways to assign 8 sections to 10 teachers.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim